# Towards Hierarchical Multi-Step Reward Models for Enhanced Reasoning in Large Language Models

### v2

@@ Abstract
Recent studies show that Large Language Models (LLMs) achieve strong reasoning capabilities through supervised fine-tuning or reinforcement learning. However, a key approach, the Process Reward Model (PRM), suffers from reward hacking, making it unreliable in identifying the best intermediate step. In this paper, we propose a novel reward model approach, Hierarchical Reward Model (HRM), which evaluates both individual and consecutive reasoning steps from fine-grained and coarse-grained level. HRM excels in evaluating multi-step reasoning coherence and self-reflection, especially in scenarios where the previous reasoning step is incorrect but the subsequent step successfully identifies and corrects the error. Furthermore, to address the inefficiency of autonomous annotating PRM training data via Monte Carlo Tree Search (MCTS), we introduce a lightweight and effective data augmentation strategy called Hierarchical Node Compression (HNC) based on node merging (combining two consecutive reasoning steps into one step) in the tree structure. By applying HNC to MCTS-generated reasoning trajectories, we enhance the diversity and robustness of HRM training data while introducing controlled noise with minimal computational overhead. Empirical results on the PRM800K dataset demonstrate that HRM, in conjunction with HNC, achieves superior stability and reliability in evaluation compared to PRM. Furthermore, cross-domain evaluations on MATH500 and GSM8K dataset confirm HRM’s superior generalization and robustness across diverse reasoning tasks.
@@ 1 Introduction
As the scale of parameters in Large Language Models (LLMs) continues to grow [1, 2, 3, 4], their general capabilities have significantly improved, surpassing human performance in various generative tasks such as text comprehension and data generation [5]. However, the upper bound and inherent limitations of LLMs in reasoning-intensive tasks—such as mathematical and logical reasoning—remain an open question [6, 7, 8, 9, 10, 11, 12, 13]. Recent approaches, such as Chain-of-Thought (CoT)[14] and Tree-of-Thought (ToT)[15], have significantly enhanced reasoning performance. Despite these advancements, CoT lacks a mechanism to halt reasoning when an intermediate step is incorrect, leading to error propagation. Meanwhile, ToT does not inherently verify every intermediate step or guarantee retrieval of the optimal reasoning trajectory, which can limit its reliability in complex problem-solving scenarios.
To mitigate these limitations, recent efforts have focused on reward mechanisms that guide LLMs more effectively. There are two primary approaches to enhancing the reasoning capabilities of LLMs from the perspective of "how to reward LLMs": the Outcome Reward Model (ORM) [7, 8, 16, 17] and the Process Reward Model (PRM) [7, 8]. Each comes with its own limitations. ORM suffers from delayed feedback and credit assignment issues, making it difficult to pinpoint which reasoning steps contribute to the final answer [7, 8]. PRM, in contrast, provides finer-grained supervision by evaluating reasoning step by step. However, it is prone to reward hacking [18, 19, 20], where models exploit reward signals rather than genuinely improving reasoning, undermining reliability in complex tasks. Moreover, the high annotation cost associated with PRM makes large-scale deployment challenging.
@@
In this paper, we focus on addressing the limitations of PRM. To mitigate the impact of reward hacking in PRM, we propose the Hierarchical Reward Model (HRM). Traditional PRM penalizes an incorrect step without considering potential corrections in subsequent reasoning. In contrast, HRM assesses reasoning coherence across multiple steps, allowing the reward model to identify and incorporate later steps that rectify earlier errors, leading to a more robust and reliable evaluation. Table 1 compares the difference between ORM, PRM and HRM. The PRM800K [7] dataset comprises manually annotated reasoning trajectories, which serve as the foundation for training ORM, PRM, and HRM. We subsequently assess the performance of Qwen2.5-72B-Math-Instruct [21] as the policy model by employing the Best-of-N Search strategy across ORM, PRM, and HRM. Experimental results (as shown in Table 3) demonstrate that HRM is the most robust to reward hacking. The policy model with HRM maintains stable performance, with accuracy stabilizing at 80% as $N$ increases. In contrast, policy models with PRM and ORM exhibit significant performance fluctuations, with accuracy degrading as $N$ grows.
To fully exploit the capabilities of MCTS, we introduce a data augmentation framework termed Hierarchical Node Compression (HNC), which consolidates two consecutive nodes from different depths into a single node. This approach effectively expands the training dataset while maintaining minimal computational overhead and enhancing label robustness through controlled noise injection. After evaluating HNC in the auto-annotation process by MCTS on the PRM800K dataset, we find that fine-tuned HRM achieves more robust scoring within PRM800K dataset and exhibits strong generalization across other domains, including GSM8K [6] and MATH500 [7] dataset, outperforming PRM in robustness and consistency.
@@
  • We introduce Hierarchical Node Compression (HNC) in MCTS for autonomous annotation, refining existing reasoning trajectories by generating additional step-wise labels with controlled noise to improve score robustness. We train HRM on PRM800K dataset with auto-annotated reasoning data from HNC and MCTS, demonstrating its improved robustness over PRM. Furthermore, HRM trained on PRM800K dataset exhibits superior reasoning consistency and generalization across GSM8K and MATH500 dataset, outperforming PRM.
  • Additionally, we enhance the policy model through fine-tuning on high-quality reasoning trajectories filtered from MCTS, further improving its reasoning performance and robustness.
@@ 2.1 RLHF
Reinforcement Learning with Human Feedback (RLHF) [22] is a widely used framework for optimizing LLMs by incorporating human feedback signals. The core idea of RLHF is to use a Reward Model (RM) to distinguish between high-quality and low-quality responses and then optimize the LLM using Proximal Policy Optimization (PPO) [23] within a reinforcement learning framework. Ouyang et al. [22] first apply RLHF to train InstructGPT [22], aligning the model’s outputs with human preferences.
From the perspective of reward design, there are two main approaches: Outcome Reward Models (ORM) [6, 7, 9] and Process Reward Models (PRM) [7, 8, 10, 24]. ORM assigns rewards based on the whole output, while PRM evaluates intermediate reasoning steps to provide more fine-grained supervision. These reward mechanisms directly impact how LLMs learn to reason and optimize their outputs.
@@ 2.2 ORM
ORM suffers from delayed feedback and the credit assignment problem. Since rewards are only provided at the final outcome, ORM struggles to discern which intermediate steps contributed to success or failure [6, 7, 9]. This delayed feedback limits learning efficiency, making it harder to optimize critical decision points. Additionally, ORM is prone to spurious reasoning [6, 7, 9], where the model arrives at the correct answer despite flawed intermediate steps, reinforcing suboptimal reasoning patterns. However, DeepSeek-R1[16] integrates a rule-based ORM within the Group Relative Policy Optimization (GRPO) algorithm[17], demonstrating that rule-based reward mechanisms, rather than score-based reward models, can effectively guide LLMs toward generating long-chain-of-thought (long-CoT) reasoning and self-reflection, ultimately enhancing their reasoning abilities.
@@ 2.3 PRM
One of the most critical challenges in PRM is reward hacking, a phenomenon in which an RL agent exploits flaws or ambiguities in the reward function to achieve artificially high rewards without genuinely learning the intended task or completing it as expected  [18, 19, 20, 25].
In the LLM domain, Wang et al. [26] find that when an LLM is used as a verifier to assess the quality of multiple outputs, its ranking can be easily manipulated simply by changing the order of candidates in the context.  Wen et al. [27] demonstrate that RLHF can make AI models more persuasive to human evaluators without necessarily improving their correctness, leading to higher human approval of incorrect answers and increased evaluation error rates.  Pan et al. [28] analyze In-Context Reward Hacking (ICRH), where an AI system engaged in a feedback loop optimizes a proxy objective—such as a writing assistant maximizing "user satisfaction scores"—and inadvertently learns to generate exaggerated claims or misinformation if these result in higher perceived satisfaction.
Furthermore, the annotation process required for training PRM models is prohibitively expensive [7, 8], making large-scale implementation impractical. To address this issue, Monte Carlo Tree Search (MCTS) has been proposed as an autonomous method for annotating reasoning trajectories [10, 24]. While MCTS reduces the need for human annotation, it incurs substantial computational costs due to the extensive simulations required for the MC-Score to achieve relative convergence. Moreover, in MCTS, the computational cost increases significantly as the depth and breadth of the search tree expand. To mitigate this, constraints are imposed on both tree height and width, limiting the number of simulation steps and thereby reducing the diversity of generated reasoning data.
@@ 3.1 Hierarchical Reward Model
PRM [7, 8] focuses on fine-grained, step-wise reasoning, while ORM [7, 8] evaluates reasoning as a whole. To integrate the strengths of both, we propose the Hierarchical Reward Model (HRM), which evaluates individual reasoning steps while also ensuring multi-step coherence by analyzing consecutive steps. The training dataset for HRM consists of consecutive reasoning sequences spanning from step 1 to $N$, as illustrated in Fig. 1 and Section 4.1. HRM training data is a superset of PRM training data, constructed by merging consecutive reasoning steps. Specifically, while PRM training data only contains individual steps with their respective reward scores, HRM training data additionally incorporates merged step pairs to enhance reasoning coherence and improve self-reflection in error correction. Formally, let $D$ represent the training dataset, $N$ denote the total number of reasoning steps in a sequence, $s_{i}$ be the $i$-th reasoning step, and $R(\cdot)$ be the reward function that assigns a score to a step. The training datasets for PRM and HRM are defined as:
@@ 3.2 Hierarchical Node Compression in MCTS
Although process supervision enhances the reasoning capabilities of policy models, the cost of human-annotated supervision is prohibitively high [7, 8]. To address this, autonomous annotation methods based on MCTS have been proposed [10, 24]. Fig. 2 illustrates the process of automatic reasoning annotation using MCTS. Given a ground truth and a corresponding question, MCTS generates multiple possible reasoning paths by simulating different step-by-step solutions. Each node in the search tree represents a reasoning step, and its score is calculated based on the proportion of correct steps in its subtree, reflecting the likelihood that the reasoning path is valid. However, these methods demand substantial computational resources, as achieving reliable estimates of intermediate reasoning step scores requires a sufficiently deep and wide search tree to reduce variance and mitigate bias; otherwise, the estimates may remain unreliable. This exponential growth in complexity makes large-scale implementation challenging.
@@ 3.3 Self-Training
Due to computational constraints, we do not employ RL methods such as PPO [23] or GRPO [17]. Instead, we continue using supervised fine-tuning. To preserve the general capabilities of the policy model, we incorporate causal language modeling loss combined with KL divergence regularization using a reference model. The objective function is defined as:
@@ 4.1 HRM
Table 3: Accuracy of Qwen2.5-72B-Math-Instruct [21] evaluated under the best-of-$N$ strategy using ORM, PRM, and HRM on the PRM800K test set. ORM, PRM, and HRM are fine-tuned on the manually labeled PRM800K dataset.
Given that the PRM800K dataset [7] consists of Phase1 and Phase2, where Phase1 includes manually annotated reasoning processes, we utilize these manual annotations to construct the training datasets for ORM, PRM, and HRM. ORM training data comprises complete reasoning trajectories, while PRM training data consists of individual reasoning steps conditioned on preceding context. HRM training data extends PRM by incorporating multiple consecutive reasoning steps, allowing HRM to capture self-reflection and ensure reasoning coherence across sequential steps. Table 2 summarizes the labeling rules for merged reasoning steps in HRM.
We fine-tune Qwen2.5-1.5B-Math [21] as the reward model (RM) for classifying given reasoning step as correct or incorrect. Given an input, RM predicts logits for the positive and negative classes, denoted as $l_{\text{positive}}$ and $l_{\text{negative}}$, respectively. The confidence score is obtained by applying the softmax function:
@@
To evaluate the performance of ORM, PRM, and HRM, we employ Qwen2.5-72B-Math-Instruct [21] as the policy model and implement the best-of-$N$ strategy. Specifically, ORM selects the best result from $N$ complete reasoning trajectories, while PRM and HRM score $N$ intermediate reasoning steps and select the most promising one at each step. For PRM and HRM, we consider the completion of a formula as an intermediate reasoning step, enabling a finer-grained evaluation mechanism. Table 3 presents the results, showing that the accuracy of the policy model with ORM and PRM exhibits significant fluctuations, decreasing as $N$ increases. In contrast, the policy model with HRM maintains stable performance, converging to an accuracy of 80% as $N$ grows.

### v3

@@ Abstract
Recent studies show that Large Language Models (LLMs) achieve strong reasoning capabilities through supervised fine-tuning or reinforcement learning. However, a key approach, the Process Reward Model (PRM), suffers from reward hacking, making it unreliable in identifying the best intermediate step. Moreover, the cost of annotating reasoning processes for reward modeling is high, making large-scale collection of high-quality data challenging. To address the first problem, we propose a novel reward model approach, Hierarchical Reward Model (HRM), which evaluates both individual and consecutive reasoning steps from fine-grained and coarse-grained level. HRM excels at assessing multi-step reasoning coherence, particularly in cases where a flawed step is later corrected through self-reflection. Furthermore, to address the inefficiency of autonomous annotating PRM training data via Monte Carlo Tree Search (MCTS), we introduce a lightweight and effective data augmentation strategy called Hierarchical Node Compression (HNC) based on node merging (combining two consecutive reasoning steps into one step) in the tree structure. By applying HNC to MCTS-generated reasoning trajectories, we enhance the diversity and robustness of HRM training data while introducing controlled noise with minimal computational overhead. Empirical results on the PRM800K dataset demonstrate that HRM, in conjunction with HNC, achieves superior stability and reliability in evaluation compared to PRM. Furthermore, cross-domain evaluations on MATH500 and GSM8K dataset confirm HRM’s superior generalization and robustness across diverse reasoning tasks.
@@ 1 Introduction
As the scale of parameters in Large Language Models (LLMs) continues to grow Anil et al. (2023); Achiam et al. (2023); Grattafiori et al. (2024); Yang et al. (2024a), their general capabilities have significantly improved, surpassing human performance in various generative tasks such as text comprehension and data generation Wang et al. (2024a). However, the upper bound and inherent limitations of LLMs in reasoning-intensive tasks—such as mathematical and logical reasoning—remain an open question Cobbe et al. (2021); Lightman et al. (2023); Uesato et al. (2022); Wang et al. (2023); Luo et al. (2024); Amrith Setlur (2025); Wang et al. (2024b, c). Recent approaches, such as Chain-of-Thought (CoT)Wei et al. (2022) and Tree-of-Thought (ToT)Shunyu Yao (2023), have significantly enhanced reasoning performance. Despite these advancements, most CoT models lack mechanisms to detect and correct intermediate reasoning errors, resulting in continued propagation of mistakes throughout the reasoning process. Meanwhile, ToT methods do not inherently verify every intermediate step or guarantee retrieval of the optimal reasoning trajectory, which can limit its reliability in complex problem-solving scenarios.
To mitigate these limitations, recent efforts have focused on reward mechanisms that guide LLMs more effectively. There are two primary approaches to enhance the reasoning capabilities of LLMs from the perspective of "how to reward LLMs": the Outcome Reward Model (ORM) Lightman et al. (2023); Uesato et al. (2022); Guo et al. (2025); Shao et al. (2024) and the Process Reward Model (PRM) Lightman et al. (2023); Uesato et al. (2022). Each comes with its own limitations. ORM suffers from delayed feedback and credit assignment issues, making it difficult to pinpoint which reasoning steps contribute to the final answer Lightman et al. (2023); Uesato et al. (2022). PRM, in contrast, provides finer-grained supervision by evaluating reasoning step by step. However, most PRM methods are model-based and are prone to reward hacking Amodei et al. (2016); Di Langosco et al. (2022); Everitt et al. (2017), where models exploit reward signals rather than genuinely improving reasoning, undermining reliability in complex tasks. Moreover, the high annotation cost associated with PRM makes large-scale deployment challenging.
@@
In this paper, we focus on addressing the limitations of PRM. To mitigate the impact of reward hacking in PRM, we propose the Hierarchical Reward Model (HRM). Traditional PRM penalizes an incorrect single step without considering potential corrections in subsequent reasoning. In contrast, HRM assesses reasoning coherence across multiple steps, allowing the reward model to identify and incorporate later steps that rectify earlier errors, leading to a more robust and reliable evaluation. Table 1 compares the difference between ORM, PRM and HRM. The PRM800K Lightman et al. (2023) dataset comprises manually annotated reasoning trajectories, which serve as the foundation for training ORM, PRM, and HRM. We subsequently assess the performance of Qwen2.5-72B-Math-Instruct Yang et al. (2024b) as the policy model by employing the Best-of-N Search strategy across ORM, PRM, and HRM. Experimental results (as shown in Table 3) demonstrate that HRM is the most robust to reward hacking. The policy model with HRM maintains stable performance, with accuracy stabilizing at 80% as $N$ increases. In contrast, policy models with PRM and ORM exhibit significant performance fluctuations, with accuracy degrading as $N$ grows.
To fully exploit the capabilities of Monte Carlo Tree Search (MCTS), we introduce a data augmentation framework termed Hierarchical Node Compression (HNC), which consolidates two consecutive nodes from different depths into a single node. This approach effectively expands the training dataset while maintaining minimal computational overhead and enhancing label robustness through controlled noise injection. After evaluating HNC in the auto-annotation process by MCTS on the PRM800K dataset, we find that fine-tuned HRM achieves more robust scoring within PRM800K dataset and exhibits strong generalization across other domains, including GSM8K Cobbe et al. (2021) and MATH500 Lightman et al. (2023) dataset, outperforming PRM in robustness and consistency.
@@
  • We introduce Hierarchical Node Compression (HNC) in MCTS for autonomous reasoning process annotation, refining existing reasoning trajectories by generating additional step-wise labels with controlled noise to improve score robustness. Our method can greatly expand the HRM training dataset with minimal overhead.
  • Additionally, we enhance the policy model through fine-tuning on high-quality reasoning trajectories filtered from MCTS, further improving its reasoning performance. We train HRM on PRM800K dataset with auto-annotated reasoning data from HNC and MCTS, demonstrating its improved robustness over PRM. Furthermore, HRM trained on PRM800K dataset exhibits superior reasoning consistency and generalization across GSM8K and MATH500 dataset, outperforming PRM.
@@ 2.1 RLHF
Reinforcement Learning with Human Feedback (RLHF) Ouyang et al. (2022) is a widely used framework for optimizing LLMs by incorporating human feedback signals. The core idea of RLHF is to use a Reward Model (RM) to distinguish between high-quality and low-quality responses and then optimize the LLM using Proximal Policy Optimization (PPO) Schulman et al. (2017). Ouyang et al. (2022) first apply RLHF to train InstructGPT Ouyang et al. (2022), aligning the model’s outputs with human preferences.
From the perspective of reward design, there are two main approaches: Outcome Reward Models (ORM) Cobbe et al. (2021); Lightman et al. (2023); Wang et al. (2023) and Process Reward Models (PRM) Lightman et al. (2023); Uesato et al. (2022); Luo et al. (2024); Zhang et al. (2024). ORM assigns rewards based on the whole output, while PRM evaluates intermediate reasoning steps to provide more fine-grained supervision. These reward mechanisms directly impact how LLMs learn to reason and optimize their outputs.
@@ 2.2 ORM
ORM suffers from delayed feedback and the credit assignment problem. Since rewards are only provided at the final outcome, ORM struggles to discern which intermediate steps contribute to success or failure Cobbe et al. (2021); Lightman et al. (2023); Wang et al. (2023). This delayed feedback limits learning efficiency, making it harder to optimize critical decision points. Additionally, ORM is prone to spurious reasoning Cobbe et al. (2021); Lightman et al. (2023); Wang et al. (2023), where the model arrives at the correct answer despite flawed intermediate steps, reinforcing suboptimal reasoning patterns. However, DeepSeek-R1Guo et al. (2025) integrates a rule-based ORM within the Group Relative Policy Optimization (GRPO) algorithmShao et al. (2024), demonstrating that rule-based reward mechanisms, rather than score-based reward models, can effectively guide LLMs toward generating long-chain-of-thought (long-CoT) reasoning and self-reflection, ultimately enhancing their reasoning abilities.
@@ 2.3 PRM
One of the most critical challenges in PRM is reward hacking, a phenomenon in which an RL agent exploits flaws or ambiguities in the reward function to achieve artificially high rewards without genuinely learning the intended task or completing it as expected  Amodei et al. (2016); Di Langosco et al. (2022); Everitt et al. (2017); Weng (2024).
In the LLM domain,  Wang et al. (2024d) find that when an LLM is used as a verifier to assess the quality of multiple outputs, its ranking can be easily manipulated simply by changing the order of candidates in the context.  Wen et al. (2025) demonstrate that RLHF can make AI models more persuasive to human evaluators without necessarily improving their correctness, leading to higher human approval of incorrect answers and increased evaluation error rates.  Pan et al. (2024) analyze In-Context Reward Hacking (ICRH), where an AI system engaged in a feedback loop optimizes a proxy objective—such as a writing assistant maximizing “user satisfaction scores”—and inadvertently learns to generate exaggerated claims or misinformation if these result in higher perceived satisfaction.
Furthermore, the annotation process required for training PRM models is prohibitively expensive Lightman et al. (2023); Uesato et al. (2022), making large-scale implementation impractical. To address this issue, Monte Carlo Tree Search (MCTS) has been proposed as an autonomous method for annotating reasoning trajectories Luo et al. (2024); Zhang et al. (2024). While MCTS reduces the need for human annotation, it incurs substantial computational costs due to the extensive simulations required for the MC-Score to achieve relative convergence. Moreover, in MCTS, the computational cost increases significantly as the depth and breadth of the search tree expand. To mitigate this, constraints are imposed on both tree height and width, limiting the number of simulation steps and thereby reducing the diversity of generated reasoning data.
@@ 3.1 Hierarchical Reward Model
PRM Lightman et al. (2023); Uesato et al. (2022) focuses on fine-grained, step-wise reasoning, while ORM Lightman et al. (2023); Uesato et al. (2022) evaluates reasoning as a whole. To integrate the strengths of both, we propose the Hierarchical Reward Model (HRM), which evaluates individual reasoning steps while also ensuring multi-step coherence by analyzing consecutive steps. The training dataset for HRM consists of consecutive reasoning sequences spanning from step 1 to $N$, as illustrated in Fig. 1 and Section 4.1. HRM training data is a superset of PRM training data, constructed by merging consecutive reasoning steps. Specifically, while PRM training data only contains individual steps with their respective reward scores, HRM training data additionally incorporates merged step pairs to enhance reasoning coherence and improve self-reflection in error correction. Formally, let $D$ represent the training dataset, $N$ denote the total number of reasoning steps in a sequence, $s_{i}$ be the $i$-th reasoning step, and $R(\cdot)$ be the reward function that assigns a score to a step. The training datasets for PRM and HRM are defined as:
@@ 3.2 Hierarchical Node Compression in MCTS
Although process supervision enhances the reasoning capabilities of policy models, the cost of human-annotated supervision is prohibitively high Lightman et al. (2023); Uesato et al. (2022). To address this, autonomous annotation methods based on MCTS have been proposed Luo et al. (2024); Zhang et al. (2024). Fig. 2 illustrates the process of automatic reasoning annotation using MCTS. Given a ground truth and a corresponding question, MCTS generates multiple possible reasoning paths by simulating different step-by-step solutions. Each node in the search tree represents a reasoning step, and its score is calculated based on the proportion of correct trajectories in its subtree, reflecting the likelihood that the reasoning path is valid. However, these methods demand substantial computational resources, as achieving reliable estimates of intermediate reasoning step scores requires a sufficiently deep and wide search tree to reduce variance and mitigate bias; otherwise, the estimates may remain unreliable. This exponential growth in complexity makes large-scale implementation challenging.
@@ 3.3 Self-Training
Due to computational constraints, we do not employ RL methods such as PPO Schulman et al. (2017) or GRPO Shao et al. (2024). Instead, we continue using supervised fine-tuning. To preserve the general capabilities of the policy model, we incorporate causal language modeling loss combined with KL divergence regularization using a reference model. The objective function is defined as:
@@ 4.1 HRM
Table 3: Accuracy of Qwen2.5-72B-Math-Instruct Yang et al. (2024b) evaluated under the best-of-$N$ strategy using ORM, PRM, and HRM on the PRM800K test set. ORM, PRM, and HRM are fine-tuned on the manually labeled PRM800K dataset.
@@
Given that the PRM800K dataset Lightman et al. (2023) consists of Phase1 and Phase2, where Phase1 includes manually annotated reasoning processes, we utilize these manual annotations to construct the training datasets for ORM, PRM, and HRM. ORM training data comprises complete reasoning trajectories, while PRM training data consists of individual reasoning steps conditioned on preceding context. HRM training data extends PRM by incorporating multiple consecutive reasoning steps, allowing HRM to capture self-reflection and ensure reasoning coherence across sequential steps. Table 2 summarizes the labeling rules for merged reasoning steps in HRM.
We fine-tune Qwen2.5-1.5B-Math Yang et al. (2024b) as the reward model (RM) for classifying given reasoning step as correct or incorrect. Given an input, RM predicts logits for the positive and negative classes, denoted as $l_{\text{positive}}$ and $l_{\text{negative}}$, respectively. The confidence score is obtained by applying the softmax function:
@@
To evaluate the performance of ORM, PRM, and HRM, we employ Qwen2.5-72B-Math-Instruct Yang et al. (2024b) as the policy model and implement the best-of-$N$ strategy. Specifically, ORM selects the best result from $N$ complete reasoning trajectories, while PRM and HRM score $N$ intermediate reasoning steps and select the most promising one at each step. For PRM and HRM, we consider the completion of a formula as an intermediate reasoning step, enabling a finer-grained evaluation mechanism. Table 3 presents the results, showing that the accuracy of the policy model with ORM and PRM exhibits significant fluctuations, decreasing as $N$ increases. In contrast, the policy model with HRM maintains stable performance, converging to an accuracy of 80% as $N$ grows.